Grade 12th passAlgebra

From the sum of 4 + 3x and 5 – 4x + 2x2 , subtract the sum of 3x2 – 5x and –x 2+ 2x + 5.

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

First we have to find out the sum of 4 + 3x and 5–4x + 2x2
= 4 + 3x + (5–4x + 2x2)
= 4 + 3x + 5–4x + 2x2
= 4 + 5 + 3x–4x + 2x2
= 9 - x + 2x2
= 2x2 –x + 9 ... [equation 1]
Then, we have to find out the sum of 3x2 –5x and–x2 + 2x + 5
= 3x2 –5x + (-x2 + 2x + 5)
= 3x2 –5x–x2 + 2x + 5
= 3x2 –x2 –5x + 2x + 5
= 2x2 –3x + 5 ... [equation 2]
Now, we have to subtract equation (2) from equation (1)
= 2x2 - x + 9–(2x2 –3x + 5)
= 2x2 –x + 9–2x2 + 3x–5
= 2x2 –2x2 –x + 3x + 9–5
= 2x + 4

Thanks